A tuning fork is used above a open tube and the 1st harmonic happens at 7 cm. Where will the 5th harmonic be heard?
Answer #1

Let the 5th harmonic will be heard at \lambda_{5} .

The wavelength of the 1st harmonic is \lambda =7 cm.

Let the length of the open tube is L.

We know for 1st harmonic

  L=\frac{1}{2}\lambda

Now we know for 5th harmonic

\lambda _{5}=\frac{2}{5}L

   or\: \: \: \: \: \: \: \lambda _{5}=\frac{2}{5}\times \left ( \frac{1}{2} \lambda \right )

or\: \: \: \: \: \: \: \lambda _{5}=\frac{2}{5}\times \left ( \frac{1}{2} \times 7 \right )

  \therefore \: \: \: \: \: \: \: \lambda _{5}=1.4\: \: \: \: \: cm

\therefore The 5th harmonic will be heard at 1.4 cm.